OpenStudy (anonymous):

if f(x)=4x^3-x^2-2x+1 and g(x)={min {f(t):0<=t<=x for 0<=x<1 and 3-x for 1

OpenStudy (anonymous):

Is this calculus?

OpenStudy (anonymous):

f(x)=4x^3-x^2-2x+1 and g(x)={min {f(t):0<=t<=x for 0<=x<1 and 3-x for 1<x<=2} then find the value of g(1/4)+g(1/2)+g(5/4)? I'm having trouble seeing what the equations are supposed to look like on here?

OpenStudy (anonymous):

2 definitions of g(x) are given !

OpenStudy (anonymous):

what are we trying to find?

OpenStudy (anonymous):

g(1/4)+g(1/2)+g(5/4)

OpenStudy (dls):

if f(x)=4x^3-x^2-2x+1 and g(x)={min {f(t):0<=t<=x for 0<=x<1 and 3-x for 1<x<=2} then find the value of g(1/4)+g(1/2)+g(5/4) \[\Large f(x)=4x^3-x^2-2x+1\] \[ g(x)=\left\{ \min \left\{ f(t):0 \leq t \leq x \right\} \right\}, 0 \leq x<1\] \[\Large g(x)=3-x , 1<x \leq 2\] ------------------------------------ We need : \[\Large g(1/4)+g(1/2)+g(5/4)\] We have 2 cases in g(x) so first we will decide which will we evaluate. if our \(\Large x>1 \) OR \(\Large x \leq 2\) and out of 1/4 , 1/2 and 5/4 only 5/4 satisfies this so.. \[\Large g(\frac{5}{4})=3-\frac{5}{4}=\frac{7}{4}\] Similarly to evaluate 1/2 and 1/4 you need to refer to the former definition and then add all 3 up!

OpenStudy (anonymous):

that is easily seen for g(5/4) how to go for g(1/2) and g(1/4)? how to evaluate the first condition?

OpenStudy (anonymous):

@DLS ?

OpenStudy (anonymous):

@ganeshie8 ?

OpenStudy (anonymous):

@Abhisar ?

OpenStudy (anonymous):

@beccaboo333 ?

OpenStudy (anonymous):

@CaseyCarns ?

OpenStudy (anonymous):

@dan815 ?

OpenStudy (anonymous):

@eric_d ?

OpenStudy (anonymous):

@Future_UsArmy_MP ?

OpenStudy (anonymous):

@hartnn ? @Hero ? @it_meh_miranda ?

OpenStudy (anonymous):

@JFraser ? @Koikkara ? @Luigi0210 ? @mathslover ? @nincompoop ? @oksuz_ ? @ParthKohli ? @qdbug @ @RosieF ? @SithsAndGiggles ? @Tushara ?

OpenStudy (anonymous):

I have no clue sry

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

anyone?

OpenStudy (jfraser):

i haven't done this kind of math in a while, sorry

OpenStudy (anonymous):

Allow me to rewrite the given functions so it's a bit clearer to see: \[f(x)=4x^3-x^2-2x+1\] \[g(x)=\begin{cases}\min\{f(t):0\le t\le x\}&\text{for }0\le x<1\\3-x&\text{for }1<x\le2\end{cases}\] You want to find \(g\left(\dfrac{1}{4}\right)+g\left(\dfrac{1}{2}\right)+g\left(\dfrac{5}{4}\right)\). \[g\left(\frac{1}{4}\right)=\min\left\{f(t):0\le t\le\frac{1}{4} \right\}\] The minimum of \(f\) over the interval \(\left[0,\dfrac{1}{4}\right]\) can be determined with some calculus, using the test for extrema. The same can be said for \(g\left(\dfrac{1}{2}\right)\).

OpenStudy (anonymous):

Hmm, the less than/equal to signs aren't showing up for me...

OpenStudy (anonymous):

\[f'(x)=12x^2-2x-2\] Critical values: \[\begin{align*}12x^2-2x-2&=0\\ 6x^2-x-1&=0\\ (3x+1)(2x-1)&=0\\ x&=-\dfrac{1}{3},~\dfrac{1}{2} \end{align*}\] The negative solution can be ignored, since \(g\) isn't defined for negative values of \(x\). Over the interval \(\left(0,\dfrac{1}{2}\right)\), \(f'(x)\) is negative, so \(f(x)\) is decreasing. Over \(\left(\dfrac{1}{2},\infty\right)\), \(f'(x)\) is positive, so \(f(x)\) is increasing. Thus \(f(x)\) has a minimum at \(x=\dfrac{1}{2}\). The minimum of \(f\) for \(\left[0,\dfrac{1}{4}\right]\) must occur at an endpoint: \[f(0)=1~~\text{and}~~f\left(\frac{1}{4}\right)=\frac{1}{2}\] So, \(g\left(\dfrac{1}{4}\right)=\min\left\{f(t):0\le t\le\dfrac{1}{4}\right\}=\dfrac{1}{2}\).

OpenStudy (anonymous):

Similarly, \(g\left(\dfrac{1}{2}\right)=\min\left\{f(t):0\le t\le\dfrac{1}{2}\right\}=\dfrac{1}{4}\), since \[f(0)=1~~\text{and}~~f\left(\frac{1}{2}\right)=\frac{1}{4}\] To recap, \[g\left(\dfrac{1}{4}\right)+g\left(\dfrac{1}{2}\right)+g\left(\dfrac{5}{4}\right)=\frac{1}{2}+\frac{1}{4}+\frac{7}{4}=\frac{5}{2}\] Kudos to @DLS for \(g\left(\dfrac{5}{4}\right)\).